# Polyhedral Complex Derivation from Piecewise Trilinear Networks

### v1

@@ 4.4 Faces and Normals
5 Implementation
We employ the HashGrid (Müller et al., 2022) for $\tau$, although any encoding using trilinear interpolation we defined can be applied. We illustrate our method in Algorithm 3 and 4, describing the details in the following sections.
5.1 Initialization
While the edge subdivision (Berzins, 2023) starts with a unit cube of eight vertices and a dozen edges, we know that the unit cube is subdivided by orthogonal planes consisting of the grid. Let an input coordinate be ${\bm{x}}\in[0,1]^{3}$, and there is a unit cube such that one of its corners is at the origin. Taking into account multi-resolution grids, we obtain the marks $m_{i}$, such that the grid planes are $x=m_{i}$, $y=m_{i}$, and $z=m_{i}$, following Algorithm 2. Notice that we carefully consider the offset $s/2$, which prevents the zero derivatives from aligning across all resolution levels (ref. Appendix A of Müller et al. (2022)). Leveraging the obtained marks, we derive the associated grid vertices and edges, which serve as the initial sets for ${\mathcal{V}}$ and ${\mathcal{E}}$, respectively.
@@ 5.1 Initialization
5.2 Optimizing Sign-vectors
5.3 Piecewise Trilinear Region
In Theorem 4.7, we need the outputs of corners ${\bm{P}}$ and ${\bm{Q}}$ in a common piecewise trilinear region; however, some corners may be outside of the region. For this, we replace all ReLU activations using the mask ${\bm{m}}$ that:
${\bm{m}}_{j}^{(i)}=(\tilde{\nu}_{j}^{(i)}({\bm{x}}_{0})>\epsilon)~{}|~{}(\tilde{\nu}_{j}^{(i)}({\bm{x}}_{7})>\epsilon),$ (20)
where ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$ are the vertices of an edge of interest, for the calculated cubic corners ${\bm{x}}_{0\dots 7}$. This implies that $\nu$ exhibits linearity, except for instances where two vertices simultaneously deviate from linearity. Moreover, as explained in Section 5.1, we can ensure that each edge resides within a shared trilinear region.
For ${\bm{P}}$ and ${\bm{Q}}$, we select the last two pairs of $i$ and $j$ such that $\gamma_{\phi(i,j)}({\bm{x}}_{0})=\gamma_{\phi(i,j)}({\bm{x}}_{7})=0$, indicating two common hypersurfaces passing two points ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$.
Algorithm 3 Polyhedral Complex Derivation
  Input: Sets of vertices ${\mathcal{V}}$ and edges ${\mathcal{E}}$ (from Section 5.1), # of layer $L$, hidden size $H$, trilinear networks $\tilde{\nu}$
  for $i=0$ to $L-1$ do
     for $j=0$ to $H-1$ do
        ${\mathcal{V}},{\mathcal{E}}=$ CurvedEdgeSubdivision$({\mathcal{V}},{\mathcal{E}},\tilde{\nu},i,j)$
     end for
  end for
  ${\mathcal{V}},{\mathcal{E}}=$ CurvedEdgeSubdivision$({\mathcal{V}},{\mathcal{E}},\tilde{\nu},L,0)$
  return ${\mathcal{V}}^{\star},{\mathcal{E}}^{\star},{\mathcal{F}}^{\star}$
Algorithm 4 Curved Edge Subdivision
  Input: A set of vertices ${\mathcal{V}}$, a set of edges ${\mathcal{E}}$, trilinear networks $\tilde{\nu}$, layer index $i$, neuron index $j$
  $d=\tilde{\nu}({\mathcal{V}};i,j)$
  ${\mathcal{V}}_{\text{new}}=\emptyset$
  for $e$ in ${\mathcal{E}}$ do
     if $d_{e_{0}}\cdot d_{e_{1}}<0$ then
        ${\bm{P}},{\bm{Q}}=$ Corners$(e)$triangleright Section 5.3
        ${\bm{x}}_{\text{new}}=$ Intersect$({\bm{P}},{\bm{Q}},e)$triangleright Theorem 4.7
        $e\leftarrow[e_{0},~{}\text{Index}({\bm{x}}_{\text{new}})]$
     end if
  end for
  ${\mathcal{V}}\leftarrow{\mathcal{V}}+{\mathcal{V}}_{\text{new}}$
  return ${\mathcal{V}}$, ${\mathcal{E}}$
6 Experiments
6.1 Hyperparameter
For ReLU neural networks, the number of layers $L$ of 3, hidden size $H$ of 16, and $\epsilon$ of 1e-4 for the sign-vectors (ref. Definition 3.4). The weight for the eikonal loss is 1e-2. For HashGrid, the resolution levels of 4, feature size of 2, base resolution $N_{\text{min}}$ of 2, and max resolution $N_{\text{max}}$ of 16 (sphere) or 32 (bunny). We use the official Python package of tinycudnn  [^4] https://github.com/NVlabs/tiny-cuda-nn for the HashGrid module.
6.2 Chamfer Distance
[…]
Figure 4 and Figure 4 show the chamfer distances for a unit sphere and the Standford bunny (Curless & Levoy, 1996). Varying the number of samples in marching cubes (Lorensen & Cline, 1987), we plot the baseline with respect to the number of generated vertices. Given that our method, Tropical, extracts vertices from the intersection points of hypersurfaces, we represent the results with a red diagonal cross.
Our approach yields nearly optimal outcomes while being parsimonious in the utilization of vertices for both scenarios. To illustrate, consider the MC achieving a comparable chamfer distance with approximately seven times more vertices and redundancy in surface representation. Specifically, the chamfer efficiency, defined as $1/\big{(}\sqrt{|{\mathcal{V}}|}\times\text{CD}\big{)}$, emphasizes this observation, revealing a ratio of 5.75 (ours) versus 4.68 (MC) as shown in Table 1.
In cases where the target surfaces exhibit planary characteristics, our method inherently holds an advantage over MC in capturing optimal vertices and faces. Nevertheless, when dealing with curved surfaces, our method may not reach the approximation achieved by MC with excessive sampling. A potential approach for our method involves breaking the curved edges to better conform to the learned surfaces. This entails allocating more vertices strategically, aiming for an improved approximation.
[…]
Sample $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$ Time $\downarrow$
MC   16     303 1364 4.21 0.00
  24     704   858 4.39 0.00
  32   1338   584 4.68 0.00
  40   2105   513 4.25 0.01
  48   3065   440 4.11 0.04
  56   4202   390 3.96 0.01
  64   5570   325 4.12 0.02
128 22612   305 2.18 0.10
196 51343   147 3.00 0.34
224 70007   161 2.34 0.62
Ours -   4491   259 5.75 5.66
Figure 2: Chamfer distances for a unit sphere varying the number of vertices.
Figure 3: Chamfer distances for the Stanford bunny varying the number of vertices.
Figure 2: Chamfer distances for a unit sphere varying the number of vertices.
Figure 3: Chamfer distances for the Stanford bunny varying the number of vertices.
Figure 4: Angular distances of the normal vectors for the Stanford bunny.
Figure 5: Effect of varying the weight of the eikonal loss on the flatness error $\Delta_{\text{flat}}$, as illustrated in Equation 24. For the plot, we conduct experiments using the unit sphere.
6.3 Angular Distance
The angular distance measures how much the normal vectors deviate from the ground truths. Analogous to the chamfer distance, we sample 200K points on the surface and calculate the normal vectors using Equation 19 in Section 4.4. The angular distance is defined as follows:
In Figure 4, the angular distances of the normal vectors for the Stanford bunny are plotted. As we expected, our approach efficiently estimates the faces using a parsimonious number of vertices compared to the sampling-based method.
6.4 Eikonal Constraint for Planarity
To assess the practical efficacy of the eikonal constraint, which enforces the planarity of hypersurfaces as outlined in Theorem 4.5, we quantify the error associated with the trilinear flat constraints presented in the proof of Theorem A.5 in Appendix. Let ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$ be two vertices consisting of an edge, while ${\bm{x}}_{1\dots 6}$ are its cubic corners. We obtain $\tilde{\nu}({\bm{x}}_{1\dots 6})$ in a piecewise trilinear region as described in Section 5.3. This evaluation is conducted as follows:
In Figure 5, we empirically validate that the eikonal loss enforces planarity in the hypersurfaces within piecewise trilinear regions, as described in Theorem 4.5. Absence of the eikonal loss results in elevated planarity errors, leading to the construction of an inaccurate mesh.
7 Limitation
In Section 4.4, the extraction of faces, including getting normal vectors and vertex sorting, gathers vertices on the same plane within a trilinear region for our visualization. To achieve this, we initially implemented a naive approach by creating a dictionary mapping regions to vertices through the iteration over the identified regions. However, we acknowledge the potential for further optimization to enhance efficiency. The process currently takes $3.23\pm 0.06$ seconds, constituting approximately 79% of the total time spent, excluding the plotting phase. Notice that we provided the detailed complexity analysis on our method in Appendix B.
8 Conclusions
In conclusion, our exploration of novel techniques in visualizing trilinear neural networks unveils intricate structures and extends mesh extraction applications beyond CPWA functions. Focused on trilinear interpolating methods, our study yields novel theoretical insights and a practical methodology for precise mesh extraction. Notably, Theorem 4.5 shows the transformation of hypersurfaces to planes within the trilinear region under the eikonal constraint. Empirical validation using chamfer distance and efficiency, and angular distance affirms the correctness and parsimony of our methodology. The correlation analysis between eikonal loss and hypersurface planarity adds depth to our understanding. Our proposed method for approximating intersecting points among three hypersurfaces, coupled with a vertex arrangement methodology, broadens the applications of our techniques. This work establishes a foundation for future research and practical applications in the evolving landscape of deep neural network analysis and real-time rendering mesh extraction.
Broader Impact and Ethical Considerations
This work aims to advance machine learning through improved mesh extraction techniques from neural surface representations. Ethical aspects align with established norms in the field, focusing on theoretical insights, analytical methods, and empirical validations for broader applications.
Acknowledgements
I would like to express my sincere appreciation to my brilliant colleagues, Sangdoo Yun, Dongyoon Han, and Injae Kim, for their contributions to this work. Their constructive feedback and guidance have been instrumental in shaping the work. The NAVER Smart Machine Learning (NSML) platform (Kim et al., 2018) had been used for experiments.
Appendix A Theoretical Proofs
Lemma A.1.
@@ Appendix A Theoretical Proofs
$=\sum_{i=0}^{2^{D-1}-1}\frac{{\bm{b}}_{D}({\bm{w}}_{D}-{\bm{a}}_{D})+{\bm{a}}_{D}({\bm{w}}_{D}+{\bm{b}}_{D})}{{\bm{a}}_{D}+{\bm{b}}_{D}}\prod_{j=1}^{D-1}\frac{(1-\textrm{B}(i)_{j}){\bm{b}}_{j}({\bm{w}}_{j}-{\bm{a}}_{j})+\textrm{B}(i)_{j}{\bm{a}}_{j}({\bm{w}}_{j}+{\bm{b}}_{j})}{{\bm{a}}_{j}+{\bm{b}}_{j}}$ (26)
@@
Proposition A.2 (Cubic Bézier curve).
@@
Lemma A.3 (Hypersurfaces in a trilinear space).
@@
Lemma A.4 (Curved edge of two hypersurfaces, restated).
@@
Notice that when $\tau({\bm{x}})_{1}+\tau({\bm{x}})_{2}+\tau({\bm{x}})_{4}=\tau({\bm{x}})_{3}+\tau({\bm{x}})_{5}+\tau({\bm{x}})_{6}=0$, at least one connecting edge can be identified. This occurs along the diagonal line ${\bm{x}}=t\cdot{\bm{1}}$, where $t\in[0,1]$. Along this line, $\tau({\bm{x}})$ represents a cubic Bézier curve, as shown in Proposition A.2 with the control points of ${\bm{P}}_{1}$ and ${\bm{P}}_{2}$ set to zeros, which is the line.
Theorem A.5 (Hypersurface and eikonal constraint, restated).
A hypersurface $\tau({\bm{x}})=0$ passing two points $\tau({\bm{x}}_{0})=\tau({\bm{x}}_{7})=0$ while $\tau({\bm{x}}_{1\dots 6})\neq 0$ satisfies the eikonal constraint $\|\nabla\tau({\bm{x}})\|_{2}=1$ for all ${\bm{x}}\in[0,1]^{3}$. Then, the hypersurface is a plane.
@@
Corollary A.6 (Affine-transformed hypersurface and eikonal constraint, restated).
@@
Similarly to Theorem A.5, we investigate the Jacobian of the eikonal constraint to be zero as follows:
@@
Since $\partial\nu_{1}/\partial\tau$ is a constant, using Lemma A.3 and replacing $\nu_{1}$ and $\tau$ with $\nu_{1}\circ\nu_{0}^{-1}$ and $\nu_{0}\circ\tau$, respectively, give us the same constaints for $\nu_{0}\big{(}\tau({\bm{x}}_{1\dots 6})\big{)}$ regardless of $\nu_{1}$. ∎
Theorem A.7 (Intersection of two hypersurfaces and a diagonal plane, restated).
@@
Then, $x,z\in[0,1]$ of the intersection point of the two hypersurfaces and a diagonal plane of $x=z$ is the solution of the following quartic equation:
@@
The solutions of the quartic equation such that $x\in[0,1]$ are the valid intersection points, and we choose one of them $x$ if there are more than one solution. Note that we can compute a batch of quartic equations using the parallel computing of the eigenvalue decomposition with Lemma A.8.
@@
We reproduce the method to find polynomial roots from companion matrix eigenvalues  (Edelman & Murakami, 1995).
Lemma A.8 (Polynomial roots from companion matrix eigenvalues, reproduced).
@@
Appendix B Complexity Analysis
We provide a detailed exposition of the complexity analysis in Section 3.4 to ensure the comprehensiveness of our work. Notice that our complexity analysis closely aligns with the approach outlined in Berzins (2023), as in their Appendix A.
@@ Appendix B Complexity Analysis
    During the initialization step (refer to Section 5.1), we determine $|{\mathcal{V}}|$ and $|{\mathcal{E}}|$ to construct a grid in $\mathbb{R}^{D}$. In the case of having $M$ HashGrid marks, the quantities are defined as $M^{D}$ and $D(M-1)M^{D-1}$, respectively. In case of the large $M$, we consider vertex and edge pruning for efficient computations inspired by a pruning strategy proposed in Berzins (2023). If, for every future neuron, denoted as $\phi(i,j)>\phi(i^{*},j^{*})$, the two vertices of an edge exhibit identical signs, the edge can be safely pruned as it will not undergo splitting. We make the assumption of that $M^{D}<|{\mathcal{V}}|$ to avoid an excessively dense grid, with a complexity of ${\mathcal{O}}(|{\mathcal{V}}|)$.
@@
In the study by Berzins (2023), a linear relationship between $|{\mathcal{V}}|$ and $|{\mathcal{E}}|$ for a mesh is empirically observed (see Figure 8a in Berzins (2023)). It is worth to note that the number of splitting edges can be effectively upper-bounded as |^$|\hat{{\mathcal{E}}}|<|{\mathcal{E}}|D/\phi(i,j)$ using Theorem 5 in (Hanin & Rolnick, 2019). Consequently, all the steps outlined in Algorithm 4 demonstrate a linear dependency on the number of vertices, denoted as ${\mathcal{O}}(|{\mathcal{V}}|)$.
To obtain faces, we sort the vertices within each region based on the Jacobian with respect to the mean of the vertices (see Algorithm 1). Considering triangularization, where each face has three vertices, the complexity remains ${\mathcal{O}}(|{\mathcal{V}}|)$, assuming the negligible cost of calculating the Jacobian, especially when performed in parallel. We recognize the possibility of further optimizations to improve efficiency for this iterative step, which we defer as part of future endeavors.
Appendix C Experimental Results
Table 2 shows the extended outcomes from Table 1, encompassing standard deviations and the results considered as ground-truth, obtained through marching cubes with 256 samples.
Table 3 shows the angular distance results depicted in Figure 4.
[…]

### v2

@@ Appendix C Algorithms
Appendix D Theoretical proofs
Lemma D.1.
@@ Appendix D Theoretical proofs
$=\sum_{i=0}^{2^{D-1}-1}\frac{{\bm{b}}_{D}({\bm{w}}_{D}-{\bm{a}}_{D})+{\bm{a}}_{D}({\bm{w}}_{D}+{\bm{b}}_{D})}{{\bm{a}}_{D}+{\bm{b}}_{D}}\cdot$
$\hskip 40.00006pt\prod_{j=1}^{D-1}\frac{(1-\textrm{B}(i)_{j}){\bm{b}}_{j}({\bm{w}}_{j}-{\bm{a}}_{j})+\textrm{B}(i)_{j}{\bm{a}}_{j}({\bm{w}}_{j}+{\bm{b}}_{j})}{{\bm{a}}_{j}+{\bm{b}}_{j}}$ (21)
@@
Proposition D.2 (Cubic Bézier curve).
@@
Lemma D.3 (Hypersurfaces in a trilinear space).
@@
Lemma D.4 (Curved edge of two hypersurfaces, restated).
@@
Notice that when $\tau({\bm{x}})_{1}+\tau({\bm{x}})_{2}+\tau({\bm{x}})_{4}=\tau({\bm{x}})_{3}+\tau({\bm{x}})_{5}+\tau({\bm{x}})_{6}=0$, at least one connecting edge can be identified. This occurs along the diagonal line ${\bm{x}}=t\cdot{\bm{1}}$, where $t\in[0,1]$. Along this line, $\tau({\bm{x}})$ represents a cubic Bézier curve, as shown in Proposition D.2 with the control points of ${\bm{P}}_{1}$ and ${\bm{P}}_{2}$ set to zeros, which is the line.
Theorem D.5 (Hypersurface and eikonal constraint, restated).
A hypersurface $\tau({\bm{x}})=0$ intersects two points $\tau({\bm{x}}_{0})=\tau({\bm{x}}_{7})=0$ while $\tau({\bm{x}}_{1\dots 6})\neq 0$ for the remaining six points. These points form a cube, with ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$ positioned on the diagonal of the cube. The hypersurface satisfies the eikonal constraint $\|\nabla\tau({\bm{x}})\|_{2}=1$ for all ${\bm{x}}\in[0,1]^{3}$. Then, the hypersurface of $\tau({\bm{x}})=0$ is a plane.
@@
Corollary D.6 (Affine-transformed hypersurface and eikonal constraint, restated).
@@
Similarly to Theorem D.5, we investigate the Jacobian of the eikonal constraint to be zero as follows:
@@
Since $\partial\nu_{1}/\partial\tau$ is a constant, using Lemma D.3 and replacing $\nu_{1}$ and $\tau$ with $\nu_{1}\circ\nu_{0}^{-1}$ and $\nu_{0}\circ\tau$, respectively, give us the same constaints for $\nu_{0}\big{(}\tau({\bm{x}}_{1\dots 6})\big{)}$ regardless of $\nu_{1}$. ∎
Theorem D.7 (Intersection of two hypersurfaces and a diagonal plane, restated).
@@
Then, $x\in[0,1]$ of the intersection point of the two hypersurfaces and a diagonal plane of $x=z$ is the solution of the following quartic equation:
@@
The solutions of the quartic equation such that $x\in[0,1]$ are the valid intersection points, and we choose one of them $x$ if there are more than one solution. Note that we can compute a batch of quartic equations using the parallel computing of the eigenvalue decomposition with Lemma D.8.
@@
We reproduce the method to find polynomial roots from companion matrix eigenvalues  [29].
Lemma D.8 (Polynomial roots from companion matrix eigenvalues, reproduced).
@@
Appendix E Complexity analysis
We provide a detailed exposition of the complexity analysis in Section 3.2 to ensure the comprehensiveness of our work. Notice that our complexity analysis closely aligns with the approach outlined in Berzins [4], as in their Appendix A.
@@ Appendix E Complexity analysis
    During the initialization step (refer to Section 5), we determine $|{\mathcal{V}}|$ and $|{\mathcal{E}}|$ to construct a grid in $\mathbb{R}^{D}$. In the case of having $M$ HashGrid marks, the quantities are defined as $M^{D}$ and $D(M-1)M^{D-1}$, respectively. In case of the large $M$, we consider vertex and edge pruning for efficient computations inspired by a pruning strategy proposed in Berzins [4]. If, for every future neuron, denoted as $\phi(i,j)>\phi(i^{*},j^{*})$, the two vertices of an edge exhibit identical signs, the edge can be safely pruned as it will not undergo splitting. We make the assumption of that $M^{D}<|{\mathcal{V}}|$ to avoid an excessively dense grid, with a complexity of ${\mathcal{O}}(|{\mathcal{V}}|)$.
@@
In the study by Berzins [4], a linear relationship between $|{\mathcal{V}}|$ and $|{\mathcal{E}}|$ for a mesh is empirically observed (see Figure 8a in Berzins [4]). It is worth to note that the number of splitting edges can be effectively upper-bounded as |^$|\hat{{\mathcal{E}}}|<|{\mathcal{E}}|D/\phi(i,j)$ using Theorem 5 in [20]. Consequently, all the steps outlined in Algorithm 2 demonstrate a linear dependency on the number of vertices, denoted as ${\mathcal{O}}(|{\mathcal{V}}|)$.
To obtain faces, we sort the vertices within each region based on the Jacobian with respect to the mean of the vertices (see Algorithm 3). Considering triangularization, where each face has three vertices, the complexity remains ${\mathcal{O}}(|{\mathcal{V}}|)$, assuming the negligible cost of calculating the Jacobian, especially when performed in parallel.
Appendix F Experimental results
Table 2 and Table 3 show the extended outcomes from the Stanford 3D Scanning dataset in Table 1. Table 4 shows the extended outcomes from the Stanford bunny of Table 1, encompassing standard deviations, including the results considered as a pseudo-ground truth, obtained through marching cubes with 256 samples. Notice that we used the more excessive 512 samples for the pseudo-ground truth of the Medium and Large models in Figure 3 to precisely measure CD after we found that 256 samples were not enough for those cases. Table 5 shows the angular distance results. Tables 2, 3, 4 and 5 use the Small model, while Table 8 uses the Large model (ref. Section 6). We conducted all experiments with an NVIDIA V100 32GB.
Table 2: Chamfer distance (CD) and chamfer efficiency (CE) for the Part A of the Stanford 3D Scanning repository [30] using the Small model as the default setting (ref. Section 6). The chamfer distance ($\times$1e-6) is evaluated using the marching cubes (MC) with 2563 grid samples as the reference ground truth for all measurements. The chamfer efficiency is defined as $100/(\sqrt{|{\mathcal{V}}|\times\text{CD}})$, providing a concise representation of the trade-off between the number of vertices and the chamfer distance. The results are averaged over three random seeds.
Bunny Dragon Happy Buddha
MC # $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$ $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$ $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$
  32   1283 4106 19.0   1416 6330 11.2   1032 6951 13.9
  64   5367 1371 13.6   6090 1936   8.5   4362 2465   9.3
128 21825   393 11.6 25236   542   7.3 18219   722   7.6
192 49569   141 14.3 57341   203   8.6 41351   276   8.8
Ours   4341   900 25.6   5104 1243 15.8   3710 1738 15.5
Table 3: Chamfer distance (CD) and chamfer efficiency (CE) for the Part B of the Stanford 3D Scanning dataset [30].
Armadillo Drill Bit Lucy
MC # $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$ $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$ $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$
  32   1438 6212 11.2   144 91959   7.5     594 6753 24.9
  64   6151 1960   8.3   873   5949 19.2   2701 2310 16.0
128 25164   605   6.6 3718   1542 17.4 11255   830 10.7
192 57306   231   7.5 8382     782 15.3 25589   335 11.7
Ours   4783 1499 13.9   694   3013 47.8   2406 1763 23.6
Table 4: Chamfer distance (CD) and chamfer efficiency (CE) for the Stanford bunny are evaluated using the marching cubes with 2563 grid samples ($\star$) as the reference ground truth. The standard deviations are provided after the $\pm$ notation, based on results obtained from three random seeds, which are used in training. Similar outcomes are observed for the remaining cases.